OpenStudy (anonymous):

I need 5 different mixed numbers that added together equal exactly 10. PLEASE HELP WILL GIVE MEDAL AND FAN TO THE FIRST PERSON TO GIVE ME A CORRECT ANSWER! Must include the 1/3 or 1/4 fractions as the fraction part of the mixed number.

OpenStudy (anonymous):

Well. \[5+(-10)+15+20+(-20)=10\]

OpenStudy (anonymous):

Dont know if that is what you're looking for.

OpenStudy (anonymous):

Without using negative numbers: \[0.5+1+1.5+2+5=10\]

OpenStudy (anonymous):

So, they all have to be mixed?

OpenStudy (anonymous):

\[0.33+1.33+2.34+3.5+2.5=10\]

OpenStudy (anonymous):

So this would be mixed numbers?\[1\frac{1}{3}+2\frac{1}{3}+3\frac{1}{3}+6\frac{1}{2}+10\frac{1}{2}=10\]

OpenStudy (danjs):

a mixed number is really like a sum already 2 1/3 = 2 + 1/3

OpenStudy (anonymous):

Yea, I got it. The reply before is an answer then :)

OpenStudy (anonymous):

\[1*1/3+2*1/3+3*1/3+12*1/4+20*1/4=10\]

OpenStudy (anonymous):

\[1\frac{1}{3}+2\frac{1}{3}+3\frac{1}{3}+12\frac{1}{4}+20\frac{1}{4}=10\]

OpenStudy (danjs):

pick a base, say 4 10 = 40/4 40/4 = (a + b + c + d + e)/4 pick some values, and make some mixed numbers

OpenStudy (danjs):

odd numbers would get you rewmainders and mixed numbers

OpenStudy (anonymous):

A mixed number has both a fraction and a whole number. 20 is a whole number @Aelia6th .

OpenStudy (danjs):

easier i think 10 = 40/4 = (a + b + c + d + e)/4 10 = 40/4 = (5 + 7 + 9 + 11 + 8) / 4 so you can use 5/4 + 7/4 + 9/4 + 11/4 + 8/4 = 10

OpenStudy (anonymous):

@Tommynaiter I know but when you add other mixed numbers and a fraction to the number 20 you get something that is far larger than 10. I think you should go here https://www.mathsisfun.com/numbers/fractions-mixed-addition.html

OpenStudy (anonymous):

@Tommynaiter Send me a message when you understand what this page is teaching you.

OpenStudy (danjs):

in terms of 1/6 ths 10 = 60/6 5 improper fractons to turn into the mixed numbers 10 = 60/6 = a/ 6 + b/6 + c/6 + d/6 + e/6 need numbers for the a though e that will total 60, and not reduce to a non mixed number a + b + c + d + e = 60 7 + 9 + 11 + 13 + 20 = 60 so, you have a sum of fractions 10 = 7/6 + 9/6 + 11/6 + 13/6 + 20/6 write in mixed form 1 1/6 + 1 1/2 + 1 5/6 + 2 1/6 + 3 1/3 those should total 10

OpenStudy (anonymous):

@DanJS